# Large Row-Constrained Supersaturated Designs for High-throughput Screening

### v1

@@ 4.1 Methods to Compare
For the sake of the comparative results in this section, we make several simplifying assumptions. First, we assume that we know the effect directions of interest. For instance, in our metallo-$\beta$-lactamase application, the goal is the inhibition of enzyme activity and so we look for a negative effect direction. Without loss of generality, we assume a positive effect direction in our simulations. Secondly, we assume a known ambient level of variability, $\sigma^{2}$, from model (1). This is a strong assumption but simplifies the analyses, particularly for OCOW and poolHiTS. Third, we assume a known background mean response, $\mu$, when there are no hits in a particular well. (We provide an alternative set of simulations in the Supplementary Material that relaxes the assumptions about $\mu$ and $\sigma^{2}$.) Finally, we assume that responses, or some standard transformation of them, can be reasonably approximated by a normal distribution.
@@ 4.1.1 One-Compound-One-Well (OCOW)
The poolHiTS procedure, using Shifted Transversal Designs (STD, Kainkaryam and Woolf, 2008), pools compounds as we suggest and also provides for row constraints. However, while its design and analysis approaches are intuitive, the STD is neither constructed using a statistically motivated criterion nor analyzed using a statistical model. In addition to this, they suffer from the following additional disadvantages compared to our proposed procedure. First, due to their cyclic construction, STDs are not available for many combinations of the numbers of compounds screened, wells and compounds per well. For example, to screen $k=96$ compounds with $n=96$ wells, one cannot construct an STD design with $c=30$ compounds per well. To obtain a higher $c$, one must decrease $n$. In our comparisons in Section 4, we had to choose $(n,k,c)$ values that STD allowed, rather than the most natural values. Real experimental screens will invariably have unique requirements, which limits the usefulness of poolHiTS. Second, the poolHiTS methodology assumes a binary response vector (that is, each well can be categorized as a “hit” if it contains an active compound, or a “miss” if it doesn’t), so in the standard numeric-response setting it is required to determine a threshold beyond which a well’s response is categorized as a “hit”.
Details can be found in Kainkaryam and Woolf (2008), but here we briefly describe the basic poolHiTS methodology. The user specifies the total number of compounds $k$ and the maximum number of compounds per well $c$, along with the anticipated maximum number of active compounds and an overall error rate. This allows the construction of a design with $a$ layers, each layer consisting of $q$ pools, where the parameters $a$ and $q$ are determined by the procedure. This results in a final STD with $n=aq$ wells, which highlights its lack of flexibility to choose the number of wells for a given $k$ and $c$. The decoding algorithm requires a binary response vector; thus, to account for uncertainty associated with such a binarization, we specify a threshold, much like that for OCOW, which we can use to label each well. We conducted some tests to establish a threshold which would best balance TPR and FPR. In the end, assuming a positive effect direction, we chose $\mu+F^{-1}(0.96)\sigma$, where $F^{-1}$ is again the inverse standard normal distribution function; that is, any wells whose response fell beyond the threshold were labeled “hits” and others assumed inert. Once the pools have been labeled, the decoding algorithm proceeds by eliminating compounds if they appear in at least $E+1$ inert wells, and tagging compounds as active if they aren’t labeled as inert and appear in at least $E+1$ positive pools. Here $E$ is the number of errors allowed in the analysis. We note that it is possible for some compounds to fall into a third inconclusive category, and in this case we label these as active since further testing would be required to eliminate them.
@@ 4.1.3 Constrained Row Screening
Our proposed approach uses a CRowS design from Section 2, constructed using 100 algorithm randomized starts of the procedure described in Section 2.1, and then analyzed using the Lasso. For the simulations, where using a profile plot to choose active effects is impractical, we use a thresholding procedure to reduce FPR, with the threshold chosen via informal testing. In particular, our analysis procedure is:
  • Center and scale the design $X$ so that each column has the same length, call it $X_{cs}$; and center response $\mathbf{y}$, call it $\mathbf{y}_{c}$;
@@ 5 Real Experiments
Metallo-$\beta$-lactamases exist in bacteria that cause nosocomial infections, including secondary bacterial infections contributing to an estimated 50% of COVID-19 deaths (Mojica et al., 2022; Chedid et al., 2021). Thus, the search for inhibitors of these enzymes is of considerable interest. To provide proof-of-concept of the CRowS approach, we performed four physical experiments for the inhibition of Verona Integron-encoded Metallo-$\beta$-lactamase 2 (VIM-2). Each of these experiments consisted of 1 known inhibitor and 30 known inert compounds, under a variety of conditions (Figure 4). The first experiment included no row constraints and an n/k ratio near 1. The other three experiments were more difficult in various ways, whether a smaller n/k ratio (Experiment 2) or row constraints (Experiments 3 and 4). The designs for each experiment were constructed using the procedure described in Section 2; they are available in the Supplementary Material. The compounds were arrayed into normalized pools implied by the designs such that each compound was at a concentration of 30 $\mu$M. We transferred the compounds into the pools using an Eppendorf epMotion 5073m liquid handing workstation. Enzymatic assays were carried out in microplates at 25C and results were determined by reading the absorbance at 495 nm in a Molecular Devices SpectraMax iD5 plate reader. The list of compounds, the four designs, and the experimental results are provided in the Supplementary Material.
@@ 6 Discussion
In this paper, we have developed a new class of supersaturated designs, motivated by applications in high-throughput screening which use multiwell plates to search for promising chemical compounds. These designs constrain the number of compounds included in each well based on logistical and biological concerns of the specific HTS setting. These designs are used as part of a method we call CRowS that use both a statistical criterion—the $UE(s^{2})$-criterion—to guide construction of the design and a statistical analysis method—the Lasso—to obtain the set of possible hit compounds. Thus, our work provides a novel and improved solution to a scientific problem as well as innovative statistical design methodology inspired by the problem. In addition, we have demonstrated an important application of supersaturated designs, including real experimental results that display the proposed methodology in action. Finally, though the proof-of-concept experiments were small, we show via simulation the use of larger SSDs than have previously been typically studied, and our methodology can scale to the hundreds or thousands of factors, opening up a rich set of computational and statistical problems for the statistical community to explore.
More specifically, in this work we have developed an efficient computational procedure to construct CRowS designs, which minimize the $UE(s^{2})$ criterion subject to $c$, the number of compounds allowed per well. We have also studied the behavior of these designs as a function of $c$, both theoretically and empirically. Theoretically, we provide initial lower bounds on the $UE(s^{2})$ criterion which can be used in some cases to verify that the globally optimal design has been obtained. Empirically, we find that when $c$ is between 40% and 45% of $k$, the constrained design is not appreciably worse than the unconstrained design. Furthermore, even severely constrained designs with $c$ as small as 5 or 10 still reliably find hit compounds in simulations. We also find that compared to the standard one-compound-one-well practice, or a non-statistical method that allows the pooling of multiple compounds in a well, CRowS detects hit compounds with substantially more reliability, while controlling the rate of false discoveries.
@@
Under the settings considered, our approach represents a clear improvement over existing HTS design and analysis methodology. However, there are several limitations to our work as well. First, though our design construction procedure is fast for the sizes of designs reported in this paper (on the order of seconds) and also for designs with $n$ and $k$ both in the hundreds (on the order of minutes), when $n$ and $k$ are both in the thousands the procedure takes on the order of days to produce a locally optimal design. This may not be a major concern because even a screening campaign studying thousands of compounds could be broken down into designs built on 96- or 384-well plates. If, however, screeners wished to experiment on, say, 1,536-well plates, they would need to be willing to wait several days for a design. Related to this are liquid handling challenges. One advantage of one-compound-one-well is that only a limited amount of liquid must be transferred from source to target plates. If instead of a single compound per well, we have $c$ compounds per well, then the time it takes to prepare the plates will increase proportionally to $c$. This, of course, must be weighed against the improved hit rates that pooling using our method could provide.
@@
Another important concern is interactions. In this work, we have assumed that compounds don’t interact either statistically or chemically. A natural question is whether statistical interactions between factors might affect the CRowS method’s ability to detect hits. In order to investigate this we repeated the simulations described in Section 4 where the response was generated using the model in (1) but the $\bm{\beta}$ included two active factors and a single active interaction. We considered both synergistic and antagonistic interactions that obey strong (interaction only appears when both parents are active) or weak (interaction only includes one active parent) heredity. In the Supplementary Material, we provide results of this simulation. We find that synergistic interactions degrade the ability of CRowS to detect hit compounds, while antagonistic interactions improve the results. We do note that Singh and Stufken (2023) has recently proposed a method to estimate interactions in SSD’s. We are currently working to adapt and apply Singh and Stufken (2023) to the present setting, so that important interactions can be accounted for and/or estimated.
On the application side, our work leads to a number of interesting questions. Though there are numerous examples of successful pooling in the literature (e.g. Devlin et al., 1996; Wilson-Lingardo et al., 1996; Taylor et al., 1997) and even in the context of commercial drug discovery (Elkin et al., 2015), there is also a concern regarding so-called “promiscuous interactions” (Feng and Shoichet, 2006), in which compounds may react to each other in unpredictable ways. In our own experimental work, before fine-tuning the process by which the designs were implemented we twice failed to achieve successful results. In the first attempt in particular, we were experimenting on a wider set of compounds than the group we have reported on in this paper. Additional experiments may clarify from the chemical side the limits of pooling in this or other contexts.
This work raises a number of theoretical questions as well. From our empirical work, it appears that row-constraint slack begins to appear at between 40% and 45% of $k$. Can this be characterized more precisely? Can we specify a threshold for $c$ beyond which we are guaranteed that we will have no row-constraint slack? Can we tighten the bounds on the designs, and perhaps develop construction methods for certain $(n,k,c)$ combinations that bypass the need for an algorithm altogether?
@@
Acknowledgements, Funding, and Disclosure Statement
This work was funded in part by Miami University’s Faculty Research Committee (BJS) and by the National Institutes of Health through grant GM128595 (RCP). We also thank Jens Mueller from Miami’s Research Computing Support for his computational assistance. The authors report there are no competing interests to declare.

### v2

@@ 4.1 Methods to Compare
For the sake of the comparative results in this section, we make several simplifying assumptions. First, we assume that we know the effect directions of interest. For instance, in our metallo-$\beta$-lactamase application, the goal is the inhibition of enzyme activity and so we look for a negative effect direction. Without loss of generality, we assume a positive effect direction in our simulations. Secondly, we assume a known ambient level of variability, $\sigma^{2}$, from model (2). This is a strong assumption but simplifies the analyses, particularly for OCOW and poolHiTS. Third, we assume a known background mean response, $\mu$, when there are no hits in a particular well. (We provide an alternative set of simulations in Section 5 of Supplementary Material A that relaxes the assumptions about $\mu$ and $\sigma^{2}$.) Finally, we assume that responses, or some standard transformation of them, can be reasonably approximated by a normal distribution.
@@ 4.1.1 One-Compound-One-Well (OCOW)
The poolHiTS procedure, using Shifted Transversal Designs (STD, Kainkaryam and Woolf, 2008), is a pooling method which also provides for row constraints. Its designs are intuitive and easy to construct, but not based upon a statistically motivated criterion. Similarly, its analysis procedure is not based upon the estimation of a statistical model. These designs also are unavailable for many $(n,k,c)$ combinations. In our comparisons in Section 4, we had to choose $(n,k,c)$ values that STD allowed, rather than the most natural values. This methodology also assumes a binary response vector (that is, that each well can be categorized as a “hit” if it contains an active compound, or a “miss” if it doesn’t), so in the standard numeric-response setting it is required to determine a threshold beyond which a well’s response is categorized as a “hit”. We conducted some tests to establish an appropriate threshold which would best balance TPR and FPR. In the end, assuming a positive effect direction, we chose $\mu+F^{-1}(0.96)\sigma$, where $F^{-1}$ is again the inverse standard normal distribution function. A description of this method is provided in Section 4 of Supplementary Material A.
@@ 4.1.3 Constrained Row Screening
Our proposed approach uses a CRowS design from Section 2, constructed using 100 algorithm randomized starts of the procedure described in Section 2.1, and then analyzed using the Lasso. For the simulations, where using a profile plot to choose active effects is impractical, we use a thresholding procedure to reduce FPR, with the threshold chosen via informal testing. In particular, our analysis procedure is the following:
  1. (a)
    Center and scale the design $X$ so that each column has the same length, call it $X_{cs}$; and center response $\mathbf{y}$, call it $\mathbf{y}_{c}$.
  2. (b)
@@ 5 Real Experiments
Metallo-$\beta$-lactamases exist in bacteria that cause nosocomial infections, including secondary bacterial infections contributing to an estimated 50% of COVID-19 deaths (Mojica et al., 2022; Chedid et al., 2021). Thus, the search for inhibitors of these enzymes is of considerable interest. To provide proof-of-concept of the CRowS approach, we performed four physical experiments for the inhibition of Verona Integron-encoded Metallo-$\beta$-lactamase 2 (VIM-2). Each of these experiments consisted of 1 known inhibitor and 30 known inert compounds, under a variety of conditions (Figure 4). The first experiment included no row constraints and an n/k ratio near 1. The other three experiments were more difficult in various ways, whether a smaller n/k ratio (Experiment 2) or row constraints (Experiments 3 and 4). The designs for each experiment were constructed using the procedure described in Section 2. The compounds were arrayed into normalized pools implied by the designs such that each compound was at a concentration of 30 $\mu$M. We transferred the compounds into the pools using an Eppendorf epMotion 5073m liquid handing workstation. Enzymatic assays were carried out in microplates at 25C and results were determined by reading the absorbance at 495 nm in a Molecular Devices SpectraMax iD5 plate reader. Additional details are in Supplementary Material A, while the list of compounds, the four designs, and the experimental results are provided in Supplementary Material D.
@@ 6 Discussion
In this paper, we have developed and studied a new class of supersaturated designs, which we call CRowS designs, motivated by applications in high-throughput screening and in particular to the search for enzymes that cause antibiotic resistance. In these designs, the number of compounds per well is constrained and we use the $UE(s^{2})$ criterion to guide their construction. We present an efficient exchange algorithm to construct them, present some initial lower bounds on the criterion value, study the effect of the row constraint upon the designs, show that they compare quite favorably to common competitors, and demonstrate that they detect a metallo-$\beta$-lactamase inhibitor in real experiments. Thus, our work presents an innovative statistical design methodology inspired by an important scientific problem, along with the suggestion to use a statistical modeling approach to the analysis of the designs. In addition, these designs are larger than have typically been studied, and our methodology can scale to the hundreds or thousands of factors.
@@
Though our approach represents a clear improvement over traditional HTS methodology, it also has several limitations. First, while our design construction procedure is fast for the sizes of designs reported in this paper, for designs with $n$ and $k$ both in the thousands (e.g. an experiment on a 1,536-well plate) may require on the order of days to produce an approximately optimal design. Also, this methodology should save on the number of required plates compared to one-compound-one-well, but may require more liquid handling since each well includes $c$ different compounds. This, of course, must be weighed against the improved hit rates that pooling using our method could provide.
@@
We also note that depending on the application, statistical or chemical interactions could be a concern. We briefly investigated a setting in which statistical interactions are sparse (this could arise, for instance, if interactions only occur between compounds that are both hits), by repeating the simulations described in Section 4 where the response was generated using the model in (2) but the $\bm{\beta}$ included two active factors and a single active interaction, half the size of the active main effects. We considered both synergistic and antagonistic interactions that obey strong (interaction only appears when both parents are active) or weak (interaction only includes one active parent) heredity. In Section 7 of Supplementary Material A, we find that synergistic interactions degrade the ability of CRowS to detect hit compounds, while antagonistic interactions improve the results. We note that Singh and Stufken (2023) has recently proposed a method to estimate interactions in SSD’s, and this could provide a possible solution to the problem of sparse interactions. There is also the possibility of “promiscuous inhibitors” (Feng and Shoichet, 2006), in which compounds may react to each other in unpredictable ways. This problem may be application-specific, or specific to particular chemical groups, and warrants additional exploration in future studies seeking to expand the number of compounds that can be simultaneously tested.
Our work raises a number of interesting follow-up questions as well. On the application side, can the methodology be used in larger-scale settings to identify potential hits? In large-scale screens, can our proposed methodology be adjusted to largely eliminate false positives in large effect environments, or be used as a secondary screening tool to eliminate false positives that are identified in the initial screen? Theoretically, can the lower bounds be tightened and the reliance on the assumption of tight constraints be relaxed? More fundamentally, is it possible to more rigorously characterize conditions that will guarantee tight row-constraints, rather than the empirical observation we’ve made that row-constraints are tight until $c$ is larger than 40%-45% of $k$? We leave these questions for future work. In general, our work opens up new applications and development for SSDs. These CRowS designs can be applied to a broad class of biological and chemical screening problems, and the size of the designs are much larger than those that have been typically studied in the SSD literature. We note that Eskridge et al. (2004) study large SSDs, and even give a nod to HTS as an application, but they don’t provide designs for $n>24$ nor do they consider row constraints or provide a real example.
@@
Acknowledgements
This work was funded in part by Miami University’s Faculty Research Committee (BJS) and by the National Institutes of Health through grant GM128595 (RCP). We also thank Jens Mueller from Miami’s Research Computing Support for his computational assistance.
Disclosure and Data Availability Statement
The authors report there are no competing interests to declare. The data underlying this article are available in its online supplementary material.